Suppose that f(x)=2/x^9. Evaluate each of the following:

f′(2) =

f′(−3) =
1
Expert's answer
2020-11-09T20:30:05-0500

"\displaystyle


 f(x) = \frac{2}{x^9} = 2x^{-9}\\

f'(x) = -18x^{-10}\\

f'(2) = -18\times 2^{-10} = \frac{-9}{512}\\


f'(-3) = -18\times (-3)^{-10} = \frac{-2}{3^8} = \frac{-2}{6561}"